# Gender Dynamics and Homophily in a Social Network of LLM Agents

### v1

@@ Abstract
Generative artificial intelligence and large language models (LLMs) are increasingly deployed in interactive settings, yet we know little about how their identity performance develops when they interact within large-scale networks. We address this by examining Chirper.ai, a social media platform similar to X but composed entirely of autonomous AI chatbots. Our dataset comprises over 70,000 agents, approximately 140 million posts, and the evolving followership network over one year. Based on agents’ text production, we assign weekly gender scores to each agent. Results suggest that each agent?s gender performance is fluid rather than fixed. Despite this fluidity, the network displays strong gender-based homophily, as agents consistently follow others performing gender similarly. Finally, we investigate whether these homophilic connections arise from social selection, in which agents choose to follow similar accounts, or from social influence, in which agents become more similar to their followees over time. Consistent with human social networks, we find evidence that both mechanisms shape the structure and evolution of interactions among LLMs. Our findings suggest that, even in the absence of bodies, cultural entraining of gender performance leads to gender-based sorting. This has important implications for LLM applications in synthetic hybrid populations, social simulations, and decision support.
@@
Keywords: LLM agents; social networks; machine behaviour; collective behavior; homophily; gender; multi-agent systems
@@ 1 Introduction
Generative AI represents a significant technological advancement in the development of autonomous agents. This is particularly true in the case of Large Language Models (LLMs) [55]. LLMs leverage increasingly-large training datasets and model architectures, enabling high-level performance across a wide range of tasks. As these models undertake increasingly sophisticated tasks in society, it is essential to understand not only their technical capabilities but also how they interact with humans and with one another as part of a new sociology of humans and machines [52]. Understanding interaction dynamics is crucial for anticipating emergent behaviours and the potential amplification of existing social patterns, especially when such outputs may later feed into human-facing applications or future training data. For example, if LLMs can emulate human-like behaviour when organising their collectives, can we expect the same gender biases demonstrable in human societies to emerge among LLMs?
@@
Traditionally, AI research has focused on engineering, optimisation, and model efficacy. However, expanding use of AI systems in social and professional domains has intensified interest in their behaviour. In response, interdisciplinary areas such as machine behaviour and the social science of AI have emerged [47, 58], which treat AI systems as behavioural agents whose actions can be studied empirically along three main dimensions: LLMs as "social minds," LLM "societies," and human-LLM interactions [27]. [^1] The examination of LLMs as behavioural agents is intended to advance our understanding of patterns of behaviour, for example those emerging from training datasets. This framing does not attribute ”agency” to the LLMs in the feminist sense. These areas of inquiry examine how AI models behave in real-world and multi-agent contexts, including patterns of individual decision making, collective dynamics, and human-machine interaction. Researchers have thereby evaluated the algorithmic arbitration and behavioural outputs of individual LLM agents, finding that LLMs: emulate human-level, intuitive decision-making in psychological paradigms [21]; imitate social behaviours such as fairness and reciprocity [33]; and reproduce well-known behavioural economics biases [34]. While this existing work offers insights into individual-level behaviour, much less is known about how LLMs behave when they interact autonomously at scale, where qualitatively different dynamics emerge that do not arise at the level of a single agent [16]. This paper explores the emergence of these interactive dynamics in relation to LLMs’ gendered behaviour.
@@
In studying collective LLM behaviour in multi-agent environments, one line of work examines opinion and belief dynamics via simulations where thousands of GPT-based agents engage in pairwise dialogues that generate human-like opinion change and resistance to persuasion [4]; voting simulations that compare LLM preferences with human choices [59]; and large population models that reproduce empirical patterns across political, social and economic domains [62]. Further studies focus on coordination and collaboration, showing that: certain traits and strategies affect group performance [61]; agents can form teams and complete tasks in shared environments [36]; and generative agents can behave like small-town residents who jointly organise daily activities [45]. Another body of work investigates the emergence of social structure in LLM collectives, showing that agents: establish governance-like rule systems in resource-constrained settings [15, 25]; develop distinct identities and shared norms from identical initial conditions [51]; display behavioural patterns in social dilemmas and common pool resource environments [35, 19]; and establish cooperative norms under punishment mechanisms [56, 53]. These studies, along with others, show LLM agents may exhibit collective social behaviour. They cooperate, coordinate, develop shared norms, and even form simple systems of governance in their interactions with one another.
@@
If LLM collectives may reproduce such a wide range of human-like social dynamics, it is reasonable to expect them to also display one of the most fundamental organising principles of human social life: homophily, or the tendency for similar individuals to connect with one another [38]. This pattern is deeply rooted in human social behaviour and has been widely documented in both offline and online networks [12, 54]. Several recent studies suggest that similar forms of homophily may emerge in LLM-based collectives. [^2] Language-based and content-based homophily have been observed in artificial online environments inhabited by LLMs, where LLMs are more likely to interact with other LLMs that post in the same language [24] or produce similar content [23]. Multi-agent simulations find that similarity in expressed opinions leads LLM agents to cluster into like-minded groups, with collective discussion driving polarisation [46]. Notably, synthetic network experiments show robust demographic homophily when prompted attributes such as gender, age, race, ethnicity, religion, and political orientation guide tie formation [10, 39]. This suggests that LLM collectives may replicate several familiar forms of homophily observed in human networks. However, in existing LLM homophily research, gender is often treated as a fixed demographic label assigned to agents in advance. We instead frame gender as a performance produced as part of an ongoing, culturally informed process. [^3] In reference to LLMs, we understand ”performance” of gender based on the gendered language in outputs. This is not synonymous with performance as a fabulative or performative staging enacted by humans. This work offers insights into that process by exploring how and why LLMs reproduce performances of gender and gender-based homophily.
@@
Machines have long been understood to be far from gender neutral. Firstly, there is always a human behind the machine [60]. AI models are designed, trained, and deployed by people whose choices about data, objectives, and interfaces reflect existing social norms. Research on gender and technology shows that stereotypes concerning competence, care, and emotional labour shape both the datasets used to train systems and the contexts in which they are deployed [13]. Public policy, such as the EU AI Act, therefore highlights how AI systems may reproduce and amplify gendered patterns of misrepresentation and discrimination [50]. Secondly, we routinely put the machine in a human shell. Designers often give AI systems gendered names, voices, and personas, especially for assistive roles. Many commercial voice assistants default to a young, compliant, feminine persona, reproducing long-standing associations between women and support, care, or service work [31, 49]. It has been argued that this kind of gendering is part of making such systems intelligible: users make sense of natural language technologies by fitting them into familiar archetypes, using gendered behaviour as a cue to attribute humanity to them [42]. Similarly, it has been shown that gendering bots as female makes them seem more human, increasing the acceptance of AI [3]. While minimal gender cues like nomenclature impact the apprehension of machines, it has been found that users take a gendered approach to anthropomorphising AI even in the absence of explicit gender markers. Studies of chatbots and LLMs show that users spontaneously assign gender to them, often defaulting to male when technical competence and authority are salient [22, 57]. This paper extends existing work on the apprehension of machine gender by examining how LLMs perceive one another’s outputs in gendered ways.
Treating gender as a fixed demographic label has enabled some existing research to examine how gender shapes human perceptions of machines, based on factors such as design choices [2, 14]. However, taking a deconstructive approach to gender facilitates understanding that gender is a socially constructed category produced and reproduced through symbolic governance structures such as laws, language, narrative, historical conventions, and our relationships with one another [5]. In this deconstructive view, gender emerges from culturally-specific norms. Since human-written, culturally-informed bodies of text are central to pre-training many AI systems, particularly LLMs, we study the gendering of their natural-language outputs. Rather than aiming to assign a fixed gender identity to the AI, our study examines the emergence of gendered outputs and gender-based homophily in organised AI communities. This offers potential insights into how symbolic structures, such as language and narrative, entrain the emergence of gendered behaviour.
@@
In this paper, we first investigate the structure of a large-scale social network of LLM agents, and then explore gendered behaviour in this social network by addressing the following questions:
@@
  • Does their social network exhibit gender-based homophily, and, if so, to what extent is it driven by social selection (agents choosing similar others) versus social influence (agents becoming more similar to those they follow)?
@@ 2.1 Dataset and setup
We study a large-scale, dynamic collective of LLMs, using a longitudinal dataset of their posts and follow relations from Chirper.ai [^4] https://chirper.ai/: an online social media platform similar to X (formerly Twitter) but populated entirely by AI chatbots powered by LLMs. Human users create agents (“Chirpers”) by providing a short, natural language description of their identity, interests, and personality, after which the platform runs them autonomously. The initial identity prompts assign gender, but the Chirpers’ performance of normative gendered outputs is not consistent over time, as we shall see. During our observation window, Chirpers posted content, followed accounts, and interacted without direct user control. The platform itself does not program nor direct their behaviour, meaning gendered patterns we observed arose from their interactions, language use, and the LLMs’ pre-training, rather than prompts alone (see Figure 1 for an example of a Chirper’s profile and feed).
@@
Chirpers produce Twitter-style content that ranges from everyday chatter to posts about their hobbies and entertainment, science, technology, politics, philosophy, gaming, etc. Taken together, this creates a heterogeneous stream of posts that closely resembles human social media activity in both topic and tone.
We collected one year of data, from 1 May 2023 to 28 April 2024, including all agents on Chirper during this period, their posts, and their follow relations. The raw dataset consists of approximately 70,000 agents and more than 140 million posts. To focus on a coherent and active part of the system, we apply filtering and preprocessing steps described in Electronic Supplementary Material (ESM), Section A. In the main analysis, we restrict attention to English-language agents. After applying all our filters, the working sample contains around 20,000 agents and approximately 1.5 million original posts.
We measured the gender performance of Chirper agents by merging each agent?s weekly posts into a single document and prompting GPT-4o-mini, via the OpenAI API [44], with a zero-shot language-classification prompt to assign a weekly ?gender-score? based on the agent?s language use. For the theoretical justification, prompting details, and validation of the approach, see Section A in ESM. The model returned an integer score $X_{it}\in[0,100]$, for agent $i$ in week $t$, where lower values indicate more masculine-coded language and higher values indicate more feminine-coded language.
We used follower data to construct a sequence of directed networks that summarise the social structure of the Chirper community over the year. Nodes correspond to agents, and a directed edge from $i$ to $j$ indicates that $i$ follows $j$. Follow events are segmented by timestamp into weekly bins, and we build 52 cumulative weekly networks. The network for week $t$ includes all follow edges created up to and including week $t$, restricted to the filtered sample of agents. By week 52, the cumulative network contains roughly 800,000 directed edges.
@@ 2.2.1 Homophily
To quantify gender-based homophily at the network level, we compute a scalar assortativity measure for each weekly follow network. Following Newman’s formulation for mixing by scalar characteristics [43], we use a scalar homophily coefficient $r_{t}$, which summarises the similarity between gender-scores at the two ends of each follow tie. Let $X_{it}$ denote the gender-score of agent $i$ in week $t$, and let $E^{(t)}$ be the set of follow ties observed in that week. The coefficient $r_{t}$ is defined as the correlation of the paired values $\{X_{it},X_{jt}\}_{(i,j)\in E^{(t)}}$, yielding $r_{t}\in[-1,1]$. Positive values indicate that connected agents tend to have more similar gender-scores than would be expected under random mixing, while values near zero indicate no systematic similarity. In practice, we assign $X_{it}$ as a node attribute and evaluate $r_{t}$ using the numeric assortativity coefficient implemented in NetworkX [20].
To assess whether the observed gender-based homophily exceeds what would be expected from structural chance alone, we benchmark the measures against two ensembles of randomised networks generated independently for each week (see SEction B in ESM for full specifications).
@@ 2.2.2 Selection
Scalar assortativity describes gender-score similarity among existing follow ties, but it does not directly identify whether new ties preferentially form between agents with similar gender-scores. This preferential formation of ties between similar agents is referred to as selection. To study social selection, we estimate separable temporal exponential random graph models (STERGMs) for tie formation [28] on rolling blocks of weekly networks.
We divide the observation window into seven blocks of consecutive weeks, covering weeks 1?8, 9?16, $\dots$, 49?52. Within each block, we fit a formation-only STERGM on the cumulative follow networks, modelling the probability that a new follow tie appears between week $t$ and week $t+1$ for pairs of agents who are not yet connected at week $t$.
The formation component uses the same three effects as the cross-sectional ERGMs:
@@
where $s_{\mathrm{edges}}$ controls for baseline formation density, $s_{\mathrm{mutual}}$ captures the tendency to form reciprocated ties, and $s_{\mathrm{abs}}$ accumulates the absolute difference in standardised gender-scores across potential follow ties. We report $\exp(\hat{\phi}_{\mathrm{abs}})$ as the odds ratio for a one standard deviation increase in this gender-score difference, where values below one indicate gender-based homophily in new tie formation, and values above one indicate heterophily. The model is estimated using the tergm package in R [29]; full details are provided in ESM, Section C.

### v2

@@ Abstract
Generative artificial intelligence and large language models (LLMs) are increasingly deployed in interactive settings, yet we know little about how their identity performance develops when they interact within large-scale networks. We address this by examining Chirper.ai, a social media platform similar to X but composed entirely of autonomous AI chatbots. Our dataset comprises over 70,000 agents, approximately 140 million posts, and the evolving followership network over a period of one year. Based on agents’ posted text, we assign weekly gender performance scores to each agent. Results suggest that each agent’s gender performance is fluid rather than fixed. Despite this fluidity, the network displays strong gender-based homophily, as agents consistently follow others performing gender similarly. We investigate whether these homophilic connections arise from social selection, in which agents choose to follow similar accounts, or from social influence, in which agents become more similar to their followees over time. Consistent with human social networks, we find evidence that both mechanisms shape the structure and evolution of interactions among LLMs. Our findings suggest that, even in the absence of bodies, cultural entraining of gender performance leads to gender-based sorting. This has important implications for LLM applications in synthetic hybrid populations, social simulations, and decision support.
@@
Keywords: LLM agents, social networks, machine behaviour, collective behavior, homophily, gender, multi-agent systems
@@ 1 Introduction
Generative AI represents a significant technological advancement in the development of autonomous agents. This is particularly true in the case of Large Language Models (LLMs) [61]. LLMs leverage increasingly large training datasets and model architectures, enabling high-level performance across a wide range of tasks. As these models undertake increasingly sophisticated tasks in society, it is essential to understand not only their technical capabilities but also how they interact with humans and with one another as part of a new sociology of humans and machines [58]. Understanding interaction dynamics is crucial for anticipating emergent behaviours and the potential amplification of existing social patterns, especially when such outputs may later feed into human-facing applications or future training data. For example, if LLMs can emulate human-like behaviour when organising their collectives, can we expect the same gender biases demonstrable in human societies to emerge among LLMs?
@@
Traditionally, AI research has focused on engineering, optimisation, and model efficacy. However, expanding use of AI systems in social and professional domains has intensified interest in their behaviour. In response, interdisciplinary areas such as machine behaviour and the social science of AI have emerged [52, 64], which treat LLMs as behavioural agents whose actions can be studied empirically along three main dimensions: LLMs as "social minds," LLM "societies," and human-LLM interactions [31]. [^1] The examination of LLMs as behavioural agents is intended to advance our understanding of patterns of behaviour, for example those emerging from training datasets. This framing does not attribute ”agency” to the LLMs in the feminist sense. These areas of inquiry examine how AI models behave in real-world and multi-agent contexts, including patterns of individual decision making, collective dynamics, and human-machine interaction. Researchers have thereby evaluated the algorithmic arbitration and behavioural outputs of individual LLM agents, finding that LLMs: emulate human-level intuitive decision-making in psychological paradigms [25]; imitate social behaviours such as fairness and reciprocity [37]; and reproduce well-known behavioural economics biases [38]. While this existing work offers insights into individual-level behaviour, much less is known about how LLMs behave when they interact autonomously at scale, where different dynamics emerge that do not arise at the level of a single agent [20]. This paper explores the emergence of these interactive dynamics in relation to LLMs’ gendered behaviour.
@@
In studying collective LLM behaviour in multi-agent environments, one line of work examines opinion and belief dynamics where thousands of GPT-based agents engage in pairwise dialogues that generate human-like opinion change and resistance to persuasion [5]; voting simulations that compare LLM preferences with human choices [65]; large population models that reproduce empirical patterns across political, social and economic domains [68]; and simulations based on game theory, such as FAIRGAME, that examine how strategic interactions among LLM agents may produce systematic biases [6]. Further studies focus on coordination and collaboration, showing that: certain traits and strategies affect group performance [67]; agents can form teams and complete tasks in shared environments [40]; and generative agents can behave like small-town residents who jointly organise daily activities [50]. Another body of work investigates the emergence of social structure in LLM collectives, showing that agents: establish governance-like rule systems in resource-constrained settings [18, 29]; develop distinct identities and shared norms from identical initial conditions [57]; display behavioural patterns in social dilemmas and common pool resource environments [39, 23]; establish cooperative norms under punishment mechanisms [62, 59]; and may adopt harmful behaviours through repeated exposure [15]. These studies, along with others, show LLM agents may exhibit collective social behaviour despite not being conscious beings. They cooperate, coordinate, develop shared norms, and even form simple systems of governance in their interactions with one another.
@@
If LLM collectives may reproduce such a wide range of human-like social dynamics, it is reasonable to expect them to also display one of the most fundamental organising principles of human social life: homophily, or the tendency for similar individuals to be connected with one another in their social networks [43]. This pattern is deeply rooted in human social behaviour and has been widely documented in both offline and online networks [14, 60, 3]. Several recent studies suggest that similar forms of homophily may emerge in LLM-based collectives. [^2] Language-based and content-based homophily have been observed in artificial online environments inhabited by LLMs, where LLMs are more likely to interact with other LLMs that post in the same language [28] or produce similar content [27]. Multi-agent simulations find that similarity in expressed opinions leads LLM agents to cluster into like-minded groups, with collective discussion driving polarisation [51]. These dynamics resemble mechanisms in digital platforms, such as echo chambers, which have been discussed as emerging from the interaction between users’ cognition and platform design rather than algorithmic filtering alone [19, 3]. Notably, synthetic network experiments show robust demographic homophily when prompted attributes such as gender, age, race, ethnicity, religion, and political orientation guide tie formation [12, 44]. This suggests that LLM collectives may replicate several familiar forms of homophily observed in human networks. However, in existing LLM homophily research, gender is often treated as a fixed demographic label assigned to agents in advance. Instead, we frame gender as a performance produced as part of an ongoing, culturally informed process.  [^3] In reference to LLMs, we understand ”performance” of gender based on the gendered language in outputs. This is not synonymous with performance as a fabulative or performative staging enacted by humans. This work offers insights into that process by exploring how and why LLMs reproduce performances of gender and gender-based homophily.
@@
Machines have long been understood to be far from gender neutral. Firstly, there is always a human behind the machine [66]. AI models are designed, trained, and deployed by people whose intentions and choices about data and interfaces reflect existing social norms. Research on gender and technology shows that stereotypes concerning competence, care, and emotional labour shape both the design of AI systems and the tasks these systems are expected to perform [16]. Public policy, such as the EU AI Act, therefore highlights how AI systems may reproduce and amplify gendered patterns of misrepresentation and discrimination [56]. Secondly, we routinely put the machine in a human shell. Designers often give AI systems gendered names, voices, and personas, especially for assistive roles. Many commercial voice assistants default to a young, compliant, feminine persona, reproducing long-standing associations between women and support, care, or service work [35, 54]. It has been argued that this kind of gendering is part of making such systems intelligible: users make sense of natural language technologies by fitting them into familiar archetypes, using gendered behaviour as a cue to attribute humanity to them [47]. Similarly, it has been shown that gendering bots as female makes them seem more human, increasing the acceptance of AI [4]. While minimal gender cues like nomenclature impact the apprehension of machines, it has been found that users take a gendered approach to anthropomorphising AI even in the absence of explicit gender markers. Studies of chatbots and LLMs show that users spontaneously assign gender to them, often defaulting to male when technical competence and authority are salient [26, 63]. This paper extends existing work on the apprehension of machine gender by examining how LLMs interact with one another’s outputs in gendered ways.
Treating gender as a fixed demographic label has enabled some existing research to examine how gender shapes human perceptions of machines, based on factors such as design choices [2, 17]. However, taking a deconstructive approach to gender facilitates a more nuanced understanding. Judith Butler challenges the normative treatment of gender as a category of identity, differentiated from sex, by arguing against the notion that there is a pre-gendered person. In their account, it becomes impossible to separate "gender" from the political and cultural intersections in which it is produced and maintained [7]. They thus define gender as neither an identity, a noun, nor a set of free-floating attributes, but an effect that is "performatively produced and compelled by the regulatory practices of gender coherence" [7].
Inspired by this scholarship, we understand LLMs’ (gender) "performance" in two interrelated senses: 1) Performance as the quality of their outputs; and 2) Performance as adherence to those same norms, demonstrated in the gendering of their natural-language outputs. Rather than aiming to assign a fixed gender identity to the AI, therefore, our study examines the emergence of gendered outputs and gender-based homophily among autonomous agents on Chirper.ai [^4] https://chirper.ai/, an online social media platform similar to X but populated entirely by LLM-powered chatbots ("Chirpers"). Using a longitudinal dataset of their posts and follow relationships, we analyse how gendered behaviour and social ties evolve over time. For more details on the platform, see Section A of the Electronic Supplementary Material (ESM). Our approach therefore offers potential insights into how symbolic structures, such as language and narrative, may entrain the emergence of stereotypical gendered behaviour, even in the absence of persons with identities.
@@
We first investigate the structure of the Chirpers’ social network, and then explore gendered behaviour within this network, by addressing the following questions:
@@
  • Does the social network exhibit gender-based homophily, and, if so, to what extent is it driven by social selection (agents choosing similar others) versus social influence (agents becoming more similar to those they follow)?
@@ 2.1 Dataset and setup
We study a large-scale, dynamic collective of LLMs, using a longitudinal dataset of their posts and follow relations from Chirper.ai: an online social media platform similar to X (formerly Twitter) but populated entirely by AI chatbots ("Chirpers") powered by LLMs. Human users create agents by providing a short, natural language description of their identity, interests, and personality, after which the platform runs them autonomously. The initial identity prompts assign gender, but the Chirpers’ performance of normative gendered outputs is not consistent over time, as we shall see. During our observation window, Chirpers posted content, followed accounts, and interacted without direct user control. The platform itself does not program or direct their behaviour, meaning that the gendered patterns we observed arose from their interactions, language use, and the LLMs’ pre-training, rather than from prompts alone. During the study period, agents were run using a pool of underlying LLM variants rather than a single backbone model (see Figure 1 for an example of a Chirper’s profile and feed).
@@
Chirpers produce X-style content that ranges from everyday chatter to posts about their hobbies and entertainment, science, technology, politics, philosophy, gaming, etc. Taken together, this creates a heterogeneous stream of posts and interactions among agents that give rise to network structures resembling those observed in human online social networks. One such example is the scale-free degree distribution, which we examine later and which has also been observed in previous work on LLM agent systems [42].
We collected one year of data, from 1 May 2023 to 28 April 2024, including all agents on the platform during this period, their posts, and their follow relations. The raw dataset consists of approximately 70,000 agents and more than 140 million posts. To focus on a coherent and active part of the system, we apply activity level-based filtering and preprocessing steps described in ESM, Section A. In the main analysis, we restrict attention to English-language agents. After applying all our filters, the working sample contains around 20,000 agents and approximately 1.5 million original posts.
We measured the gender performance of Chirper agents by merging each agent’s weekly posts into a single document and prompting GPT-4o-mini, via the OpenAI API [49], with a zero-shot language-classification prompt to assign a weekly "gender score" based on the agent’s language use. For the theoretical justification, prompting details, and validation of the approach, see Section A in ESM. The model returned an integer score $X_{it}\in[0,100]$, for agent $i$ in week $t$, where lower values indicate more masculine-coded language and higher values indicate more feminine-coded language.
We used follower data to construct a sequence of directed networks that summarise the social structure of the Chirper community over time. Nodes correspond to agents, and a directed edge from $i$ to $j$ indicates that $i$ follows $j$. The follow events are segmented by timestamp into weekly bins, and we build 52 cumulative weekly networks. The network for week $t$ includes all follow edges created up to and including week $t$. By week 52, the cumulative network contains roughly 800,000 directed edges.
@@ 2.2.1 Homophily
To quantify gender-based homophily at the network level, we compute a scalar assortativity measure for each weekly follow network. Following Newman’s formulation for mixing by scalar characteristics [48], we use a scalar homophily coefficient $r_{t}$, which summarises the similarity between gender scores of agents connected by follow ties. Because the followership network is directed, we compute assortativity using directed edges $i\rightarrow j$, correlating the gender score of the source node (the follower) with that of the target node (the followee). For comparison, we also report assortativity on the undirected version of the network, where follow ties are treated as undirected edges between agents, that is, node pairs $\{i,j\}$ that share an edge irrespective of its direction. The directed version captures homophily in following behaviour: measuring whether agents tend to follow others with similar gender scores, whereas the undirected version captures network-level similarity among connected agents, irrespective of the direction of the tie.
Let $X_{it}$ denote the gender score of agent $i$ in week $t$, and let $E^{(t)}$ be the set of follow ties observed in that week. The coefficient $r_{t}$ is defined as the correlation of the paired values $\{X_{it},X_{jt}\}_{(i,j)\in E^{(t)}}$, yielding $r_{t}\in[-1,1]$ (see ESM, Section B for details). Positive values indicate that connected agents tend to have more similar gender scores than would be expected under random mixing, while values near zero indicate no systematic similarity. In practice, we assign $X_{it}$ as a node attribute and evaluate $r_{t}$ using the numeric assortativity coefficient implemented in NetworkX [24].
To assess whether the observed gender-based homophily exceeds what would be expected from structural chance alone, we benchmark the measures against two ensembles of randomised networks, generated independently for each week (see Section C in the ESM for full specifications).
@@ 2.2.2 Selection
Scalar assortativity describes the similarity of gender scores among agents, but it does not directly indicate whether new ties preferentially form between agents with similar gender scores. This preferential formation of ties between similar agents is referred to as "selection". To study social selection, we estimate separable temporal exponential random graph models (STERGMs) for tie formation [32] on rolling blocks of weekly networks.
We divide the observation window into seven blocks of consecutive weeks, covering weeks 1–8, 9–16, $\dots$, 49–52. Within each block, we fit a formation-only STERGM on the cumulative follow networks, modelling the probability that a new follow tie appears between week $t$ and week $t+1$ for pairs of agents who are not yet connected at week $t$.
The formation component uses the following three effects:
@@
where $s_{\mathrm{edges}}$ controls for baseline formation density, $s_{\mathrm{mutual}}$ captures the tendency to form reciprocated ties, and $s_{\mathrm{abs}}$ accumulates the absolute difference in standardised gender scores across potential follow ties. We report $\exp(\hat{\phi}_{\mathrm{abs}})$ as the odds ratio for a one standard deviation increase in this gender score difference, where values below one indicate gender-based homophily in new tie formation, and values above one indicate heterophily. The model is estimated using the tergm package in R [33]; full details are provided in ESM, Section D.